OpenStudy (anonymous):

A right circular cone has a lateral area of 1,280π ft.^2 and a base area of 512π ft.^2 Find the volume.

OpenStudy (anonymous):

What are the equations for the LA and BA of a cone?

OpenStudy (pawanyadav):

Do you know what is base area?

OpenStudy (anonymous):

is the base area pi r ^2? and the lateral is pi rl

OpenStudy (pawanyadav):

Yes now put Base area =512π And lateral area=1280π And find r,l?

OpenStudy (pawanyadav):

Can you do this..

OpenStudy (anonymous):

the r = 22.63 am i right?

OpenStudy (pawanyadav):

Don't convert it to decimal keep it in square root form so r=√(512)

OpenStudy (pawanyadav):

Now what is l

OpenStudy (anonymous):

l = 40 square root 2?

OpenStudy (pawanyadav):

Correct now you know r=16√2 l=40√2 ,try to find altitude

OpenStudy (anonymous):

i'm sorry but what is the formula of altitude in cone?

OpenStudy (pawanyadav):

h^2=l^2-r^2

OpenStudy (anonymous):

\[8\sqrt{42}\]

OpenStudy (anonymous):

?

OpenStudy (pawanyadav):

That should be correct

OpenStudy (anonymous):

and then what will be next step?

OpenStudy (pawanyadav):

Now put this in 1/3πr^2h

OpenStudy (anonymous):

and the answer is 27,797.98. right?

OpenStudy (pawanyadav):

Yes ,I got this too

OpenStudy (anonymous):

yay! thank you again and again. :))

OpenStudy (pawanyadav):

~27798

OpenStudy (pawanyadav):

Welcome

OpenStudy (pawanyadav):

This one

OpenStudy (anonymous):

yay! thank you again and again. :))